Consider a sample with data values of 26, 24, 23, 16, 32, 33, 28, and 24. Compute the range, interquartile range, variance, and standard deviation (to a maximum of 2 decimals, if decimals are necessary. Use Excel's =quartile.exc function).
Answer #1

So , Range = Highest value - lowest value = 33-16 = 17

IQR = Q3 - Q1

Now using =quartile.exc function , and putting 1 and 3 in "quart" we get Q1=23.25 and Q3 = 31

So, IQR = 31-23.25 = 7.75

Using excels "=VAR.S " function we get , variance = 29.36

and standard deviation